A parallel plate capacitor with plate separation of 4.0 cm has a plate area of 4.0 × 10-2 m2. What is the capacitance of this capacitor with air between these plates?

a) 8.9 × 10-11 F
b) 8.9 × 10-15 F
c) 8.9 × 10-12 F
d) 8.9 × 10-14 F
e) 8.9 × 10-13 F
Answer #1

8.9 X 10^-12
Answer #2
8.9 × 10-12 F
Answer #3
C = 8.85*10^-12*4*10^-2/0.04 =8.9 *10^-12 F =8.9 pF

option C